OpenStudy (anonymous):

Solve 6t = 78. Solve 6t = 78. @Mathematics

OpenStudy (anonymous):

hahahahaha

OpenStudy (anonymous):

is 72

OpenStudy (anonymous):

just divide the RHS by 6 and see wt u get

OpenStudy (anonymous):

@wlbrt316 : You sure you need answer to that one? 6t=72 multiply by 1/6 on both sides 6t/6 = 72/6 hence t = 12

OpenStudy (anonymous):

is 13

OpenStudy (anonymous):

6t=78

OpenStudy (anonymous):

great wlbrt keep it up do it by urself don't rely on others

OpenStudy (anonymous):

wlbrt316, you said it is 72